35x+7y=32
-5x=y+20

Answer by mananth(16946)

35x+7y=32
7y=-35x+32
y=-35x/7 +32/7............1
-5x=y+20
multiply by 7
-35x=7y+140
7y=-35x-140
y=-35x/7 -20
the slope is same for both lines
and the y intercepts are different
so the lines are parallel and have no solution
35 x + 7 y = 32 .............1
-5 x -1 y = 20 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 7
35 x 7 y = 32
-35 x -7 y = 140
Add the two equations
0 x = 172
Not possible